OpenStudy (anonymous):

equation for a hyperbola with vertices: (23,0) and (-23,0), conjugate axis of length 4 units

OpenStudy (amistre64):

if we allow for an y/x slope that corresponds to b/a .... x = 23, y is equal to ... i believe half of the axis length: so 2

OpenStudy (amistre64):

since the verts given ride the x axis, x is positive \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] \[\frac{x^2}{23^2}-\frac{y^2}{2^2}=1\]